OpenStudy (anonymous):

If the side length of a square pyramid is tripled and the slant height is divided by 5. What would be the formula to find the modified surface area

OpenStudy (anonymous):

Finding the actual formula, gimma a sec

OpenStudy (anonymous):

SA = s^2 + 2 * s * l

OpenStudy (anonymous):

so this would be the new formula: \[SA =\frac{ s^{2} }{ 5 } + 2\frac{ s }{ 5 } * l(3)\]